# FaVoR: Features via Voxel Rendering for Camera Relocalization

### v1

@@ 4 Results
Dataset Feature Median Pose Error (cm, deg) $\downarrow$ Avg. Inliers per Image $\uparrow$
Extractor Iter 1 Iter 2 Iter 3 Iter 1 Iter 2 Iter 3
@@
We evaluated our system for camera relocalization in both indoor and outdoor environments using two well-established benchmarks: the 7-Scenes [40] and Cambridge Landmarks [17] datasets. We provide implementation details and information about our baseline comparisons in Section 4.1. In Section 4.2, we evaluate our method against existing state-of-the-art (SOTA) relocalization systems that use different scene representation paradigms. We report on the memory footprint and discuss the computational requirements of FaVoR compared to other systems in Section 4.3. Lastly, in Section 4.4, we examine the rendering capabilities of FaVoR as a function of the viewing angle.
@@ 4.1 Implementation Details and Baselines
Our implementation is based on PyTorch, building on the framework provided by [42]. We use Alike [50] and SuperPoint [11] as feature extractors. We test descriptors with 64 (Alike-t), 94 (Alike-s), 128 (Alike-n), 128 (Alike-l), and 256 (SuperPoint) channels. Training is carried out for 2,000 epochs per voxel (landmark) using 1,024 rays per epoch, with the Adam optimizer [18] and a variable learning rate per subvoxel as in [42]. We examine the tradeoff between model size, rendering quality, expressed by the signal-to-noise ratio (PSNR), and grid resolution (details in the supplementary material), and chose a grid resolution of $3\times 3\times 3$ to balance memory usage and rendering performance. Training a voxel on a $7\times 7$ pixels descriptor patch takes about 10 seconds on the NVIDIA GeForce RTX 4060 laptop GPU, rendering a single descriptor takes about 1 ms. However, FaVoR is not currently optimized for fast runtime operations. Since the voxels are independent of each other, the entire training and inference pipeline can be parallelized to significantly improve performance.
We compare FaVoR with similar SFR methods that render descriptors from unseen views: FQN [14], CROSSFIRE [28], and NeRF-loc [21], which closely align with our work in terms of scope, despite the different methodologies and requirements, as discussed in Section 2. Contrary to FQN [14] and CROSSFIRE [28] that utilize a large set of points provided by SfM and dense feature matching, respectively, or NeRF-loc [21] which uses 1,024 3D points for matching each query image, we use a smaller set of points similarly to [10]. Additionally, while [21, 28] require a fully trained neural rendering model to work, FaVoR needs only a sequence of frames to train the sparse voxel representation of the landmarks. Moreover, FaVoR works with out-of-the-box feature extractor methods, without requiring a custom, scene-dependent network as in [28]. With the training of FaVoR using descriptors with different numbers of channels, we demonstrate that our method is scalable to different descriptor sizes, unlike [14]. Also, [14] iterates the Render+PnP-RANSAC multiple times, we use only three iterations, similarly to [28]. For comprehensiveness, we also report the localization results of image-based [17, 6, 39, 29] and structure-based [46, 5, 3] camera relocalization methods, as well as hybrid methods (see Section 2 for more details). For the HMs category, we report the results from HLoc [35], based on SuperPoint [11] (SP) and SuperGlue [36] (SG), which uses a SfM point cloud for relocalization. We also include results for SceneSqueezer [48] which requires less memory but at the cost of higher localization error.
@@ 4.2 Relocalization Evaluation
Following the setup described in Section 4.1, in Table 1, we observe an increase in the number of inliers, accompanied by the decrease of the rotation and translation errors as the camera pose approaches the (true) camera pose. This observation, suggests that FaVoR faithfully produces descriptors not only from poses seen during training but also from new views.
@@
To evaluate the impact of poor initialization on localization, we used the first pose per scene from the 7-Scenes [40] dataset as a uniform prior for all images, instead of DenseVLAD [43]. This simulates increasing noise as the query pose gets further from the initial pose. As shown in Table 2 (more results in the supplementary), despite imprecise priors, the Render+PnP-RANSAC iterations reliably deliver accurate localization, demonstrating the robustness of our method to initial pose variations.
@@
Indoor: 7-Scenes. The 7-Scenes[40] dataset includes a sequence of frames recorded with a Kinect RGB-D camera, representing seven different indoor environments. The recordings present motion blur, small baselines between the frames, and occlusion, typical elements for real-life indoor robotics applications. We use the full image size of 640$\times$480 pixels for training and testing. For the evaluation, we used the poses provided by [4] as ground truth. In Table 3, we report the median translation and rotation errors (in centimeters and degrees) for state-of-the-art pose relocalization methods divided into categories according to the data representation paradigm. Our system outperforms the SOTA SFR methods by 39% and 69% on the translational and rotational errors, respectively. FaVoR also delivers better results than the SBMs, IBMs, and HLoc [35] approaches. In particular, for the indoor dataset, we report the results from the implementation repository of HLoc [35], which are obtained using the depth maps provided by [5]. We limited the number of landmarks for the indoor environments to 1,500.
@@
Outdoor: Cambridge Landmarks. The Cambridge Landmarks [17] dataset includes images from large outdoor scenes that are 875 to 5,600 $\text{m}^{2}$ in size, with significant view and appearance changes. We reduce the size of the images to 512$\times$288 pixels for training and testing, as Alike [50] performs faster with lower resolutions. For outdoor testing, we optimized 5,000 voxels for the Shop and Hospital scenes, and 10,000 for the others, based on scene size. As shown in Table 4, similarly to the indoor environment, our system outperforms the SB and IB methods, whereas the hybrid methods show their efficacy for large scenarios, achieving the best performance. For the evaluation of HLoc [35], we report the results from LIMAP [22], which, unlike the original HLoc, account for the camera distortion model during landmark reprojection, providing the best results for outdoor environments. FaVoR provides the second-best camera pose estimation results in the SFR category. In particular, FaVoR yields better results than FQN [14] and CROSSFIRE [28] in all the scenarios, while NeRF-loc [21] consistently provides an error of 1-3 centimeters lower than FaVoR. One of the main reasons behind the lower performance compared to NeRF-loc [21] lies in the inaccurate estimate of the position of the landmarks computed during the triangulation step. Unlike NeRF-loc [21], which utilizes fully trained NeRF models and RGB-D images for accurate 3D scene rendering, we adopt a sparse visual localization front-end that triangulates landmarks using tracks, which can introduce inaccuracies. The pose estimation errors observed with FaVoR in the outdoor dataset may result from insufficient viewpoint variation, which makes accurate triangulation challenging given the large scene depth. However, despite the lower accuracy of FaVoR for outdoor scenes compared to NeRF-loc [21], our method with Alike, as analyzed in Section 4.3, uses less memory than [21] and does not need a complete neural rendering of the scene.
@@ 4.3 Memory Usage and Training Complexity
Table 5: Scene representation size comparison. Comparison between the size of the largest checkpoints obtained with FaVoR. The Chess [40] and King’s College [17] scenes results in larger checkpoints compared to the other scenes. For HLoc [35], we report the memory used to store the features only in the two scenes.
In Table 5, we compare the checkpoint memory requirements of FaVoR to represent a scene for both datasets against other methods. FaVoR uses more memory than FQN [14] because it represents each landmark independently. However, this enables our representation to scale effectively, rendering high-dimensional descriptors in large scenes and consistently outperforming FQN [14]. On the other hand, FaVoR uses less than half the memory required by CROSSFIRE [28]. However, for large outdoor scenarios, our method requires more memory. Despite the larger memory footprint, FaVoR delivers consistently better localization performance than CROSSFIRE [28]. FaVoR also requires less memory compared to NeRF-loc [21] for indoor scenes and outdoor with Alike descriptors. Our method shows its potential especially when compared to hybrid methods. For instance, FaVoR uses more than 200 times less memory than HLoc [35] with SP+SG [11, 36] in indoor scenes. Methods like SceneSqueezer [48] significantly reduce the memory requirements (0.3 MB) for HBs, but with significantly increased localization error. Image-based methods generally require less memory than structure-based ones but with the caveat that they provide less accurate pose estimates in general. Advantages of our method include its ability to train a voxel from feature tracks, making it suitable for robotics localization pipelines.
@@ 4.4 Rendering Capabilities
When extracting feature points from an image (i.e., keypoints and associated descriptors), the descriptors must be largely view-invariant to ensure robust matching capabilities from different viewpoints. To evaluate the view-invariance of feature descriptors, we extract dense descriptor maps from a sequence of images taken from different angles. For each image, we use Alike-l [50] to extract features from a target image and compute similarity scores between each feature and the dense maps. The same process is applied to descriptors rendered by FaVoR, but using the ground truth pose for rendering.
Figure 3 shows the median values of the top thirty similarity scores per image at different query angles, for Alike-l and FaVoR trained to render Alike-l features. FaVoR maintains almost constant values, indicating that the descriptors are rendered faithfully from unseen views close to the ones observed during training. In contrast, there is a noticeable degradation in the median similarity scores when attempting to match features directly extracted by Alike-l, with a significant drop occurring beyond ±$\pm 30$ degrees, as highlighted in the plot. Figure 3 demonstrates the effectiveness of FaVoR in maintaining descriptor rendering fidelity across varying viewpoints.
@@ 5 Conclusion
In this paper, we presented FaVoR, a novel, globally sparse and locally dense vision-based localization method. Our approach involves tracking features over a sequence of RGB images with known poses and learning a 3D descriptor representation that models variations in appearance according to observation viewpoints. Our method is also suitable for integration with existing track-based localization systems for robotics.

### v4

@@ 4 Results
Dataset Feature Extractor Median Pose Error (cm, deg) $\downarrow$ Avg. Inliers per Image $\uparrow$
Iter 1 Iter 2 Iter 3 Iter 1 Iter 2 Iter 3
@@
We evaluate our system for camera relocalization in both indoor and outdoor environments using two well-established benchmarks: the 7-Scenes [40] and Cambridge Landmarks [17] datasets. We provide implementation details and information about our baseline comparisons in Section 4.1. In Section 4.2, we evaluate our method against existing state-of-the-art (SOTA) relocalization systems that use different scene representation paradigms. We report on the memory footprint and discuss the computational requirements of FaVoR compared to other systems in Section 4.3. Lastly, in Section 4.4, we examine the rendering capabilities of FaVoR as a function of the viewing angle.
@@ 4.1 Implementation Details and Baselines
Our implementation is based on PyTorch, building on the framework provided by [42]. We use Alike [50] and SuperPoint [11] as feature extractors. We test descriptors with 64 (Alike-t), 94 (Alike-s), 128 (Alike-n), 128 (Alike-l), and 256 (SuperPoint) channels. Training is carried out for 2,000 epochs per voxel (landmark) using 1,024 rays per epoch, with the Adam optimizer [18] and a variable learning rate per subvoxel as in [42]. We examine the tradeoff between model size, rendering quality, expressed by the signal-to-noise ratio (PSNR), and grid resolution (details in the supplementary material), and chose a grid resolution of $3\times 3\times 3$ to balance memory usage and rendering performance. Training a voxel on a $7\times 7$ pixel descriptor patch takes about 10 seconds on an NVIDIA GeForce RTX 4060 laptop GPU, rendering a single descriptor takes about 1 ms. However, FaVoR is not currently optimized for fast runtime operations. Since the voxels are independent of each other, the entire training and inference pipeline can be parallelized to significantly improve performance.
We compare FaVoR with similar SFR methods that render descriptors from unseen views: FQN [14], CROSSFIRE [28], and NeRF-loc [21], which closely align with our work in terms of scope, despite the different methodologies and requirements (see Section 2). FQN [14] and CROSSFIRE [28] rely on large sets of points, using SfM and dense feature matching, respectively. NeRF-loc [21] matches each query image with 1,024 3D points. In contrast, our approach uses a smaller set of points. Additionally, while [21, 28] require a fully trained neural rendering model to work, FaVoR needs only a sequence of frames to train the sparse voxel representation of the landmarks. Moreover, FaVoR works with out-of-the-box feature extraction methods, without requiring a custom, scene-dependent network as in [28]. By training FaVoR using descriptors with different numbers of channels, we demonstrate that our method is scalable to different descriptor sizes, unlike [14]. Also, while [14] iterates the Render+PnP-RANSAC multiple times, we use only three iterations, similarly to [28]. For comprehensiveness, we report the localization results of image-based [17, 6, 39, 29] and structure-based [46, 5, 3] camera relocalization methods, as well as hybrid methods (see Section 2 for more details). For the HM category, we report the results from HLoc [35], based on SuperPoint [11] (SP) and SuperGlue [36] (SG), which uses a SfM point cloud for relocalization. We also include results for SceneSqueezer [48] which requires less memory but at the cost of greater localization error.
@@ 4.2 Relocalization Evaluation
Following the setup described in Section 4.1, we observe in Table 1 an increase in the number of inliers, accompanied by a decrease in the rotation and translation errors as the camera pose approaches the (true) pose. This observation, suggests that FaVoR faithfully produces descriptors from poses seen during training as well as for new views.
@@
To evaluate the impact of poor initialization on localization, we used the first pose per scene from the 7-Scenes dataset [40] as a uniform prior for all images, instead of DenseVLAD [43] pose priors.This simulates increasing noise as the query pose moves further from the initial pose. As shown in Table 2 (more results in the supplementary), despite imprecise priors, the Render+PnP-RANSAC iterations reliably yield accurate localization, demonstrating the robustness of our method to initial pose variations.
@@
Indoor: 7-Scenes. The 7-Scenes[40] dataset includes a sequence of frames recorded with a Kinect RGB-D camera, of seven different indoor environments. The data presents motion blur, small baselines between the frames, and occlusion, all typical elements for real-life indoor robotics applications. We use the full image size of 640$\times$480 pixels for training and testing. For evaluation, we used the poses provided by [4] as ground truth. In Table 3, we report the median translation and rotation errors (in centimetres and degrees) for state-of-the-art (SOTA) pose relocalization methods, divided into categories according to the data representation paradigm. Our system outperforms the SOTA SFR methods by 39% and 69% for the translational and rotational errors, respectively. FaVoR also yields better results than the SBMs, IBMs, and HLoc [35] approaches. In particular, for the indoor dataset, we report the results from the implementation repository of HLoc [35], which are obtained using the depth maps provided by [5]. We limited the number of landmarks for the indoor environments to 1,500.
@@
Outdoor: Cambridge Landmarks. The Cambridge Landmarks [17] dataset includes images from large outdoor scenes that are 875 to 5,600 $\text{m}^{2}$ in size, with significant view and appearance changes. We reduce the size of the images to 512$\times$288 pixels for training and testing, since Alike [50] is faster when using lower-resolutions. For testing on the outdoor data set, we optimized 5,000 voxels for the Shop and Hospital scenes, and 10,000 for the others, based on scene size. As shown in Table 4, similar to the indoor environment results, our system outperforms the SB and IB methods, with the hybrid methods achieving the best performance in large scenarios. For the evaluation of HLoc [35], we report the results from LIMAP [22], which, unlike the original HLoc, account for the camera lens distortion model when reprojecting the landmarks, providing the best results for outdoor environments. FaVoR produces the second-best camera pose estimation results in the SFR category. In particular, FaVoR yields better results than FQN [14] and CROSSFIRE [28] for all of the scenarios, while NeRF-loc [21] consistently provides an error of 1-3 centimeters lower than FaVoR. One of the main reasons for the reduced performance relative to NeRF-loc [21] is the inaccurate estimate of the position of the landmarks computed during the triangulation step. Unlike NeRF-loc [21], which utilizes a fully trained NeRF model and RGB-D images for accurate 3D scene rendering, we adopt a sparse visual localization front-end that triangulates landmarks using tracks, which could introduce inaccuracies. The pose estimation errors observed with FaVoR in the outdoor dataset may result from insufficient viewpoint variation, which makes accurate triangulation challenging given the large scene depth. However, despite the lower accuracy of FaVoR for outdoor scenes, our method with Alike, as analyzed in Section 4.3, uses less memory than [21] and does not require a complete neural rendering of the scene.
@@ 4.3 Memory Usage and Training Complexity
Table 5: Scene representation size comparison. Comparison between the size of the largest checkpoints obtained with FaVoR. The Chess [40] and King’s College [17] scenes result in larger checkpoints than the other scenes. For HLoc [35], we report the memory used to store the features only in the two scenes.
In Table 5, we compare the checkpoint memory requirements of FaVoR for both datasets against other methods. FaVoR uses more memory than FQN [14] as our method represents each landmark independently. However, this enables our representation to scale effectively, rendering high-dimensional descriptors in large scenes and consistently outperforming FQN [14]. On the other hand, FaVoR uses less than half the memory required by CROSSFIRE [28]. However, for large outdoor scenes, our method does require more memory. Despite the larger memory footprint, FaVoR delivers consistently better localization performance than CROSSFIRE [28]. FaVoR Alike-l also requires less memory compared to NeRF-loc [21] for indoor and outdoor scenes.
Our approach also performs favorably compared to hybrid methods. For instance, FaVoR uses over 200 times less memory than HLoc [35] with SP+SG [11, 36] in indoor scenes. Methods like SceneSqueezer [48] significantly reduce the memory requirements (0.3 MB) for HBs, but at the cost of increased localization error. Image-based methods generally require less memory than structure-based ones but with the caveat that they provide less accurate pose estimates in general. Advantages of our method include its ability to train a voxel from feature tracks, making it suitable for robotics localization pipelines.
@@ 4.4 Rendering Capabilities
When extracting feature points from an image (i.e., keypoints and associated descriptors), the descriptors must be largely view-invariant to ensure robust matching from different viewpoints. To evaluate the view-invariance of feature descriptors, we extract dense descriptor maps from a sequence of images taken from different angles. For each image, we use Alike-l [50] to extract features from a target image and compute similarity scores between each feature and the dense maps. The same process is applied to descriptors rendered by FaVoR, but using the ground truth pose for rendering.
Figure 3 shows the median values of the top thirty similarity scores per image at different query angles, for Alike-l and FaVoR trained to render Alike-l features. FaVoR maintains an almost constant value, indicating that the descriptors are rendered faithfully from unseen views close to the ones observed during training. In contrast, there is a noticeable degradation in the median similarity scores when attempting to match features directly extracted by Alike-l, with a significant drop occurring beyond ±$\pm 30$ degrees. Figure 3 demonstrates the effectiveness of FaVoR in maintaining descriptor rendering fidelity across varying viewpoints.
@@ 5 Conclusion
In this paper, we presented FaVoR, a novel, globally sparse and locally dense vision-based camera relocalization method. Our approach involves tracking features over a sequence of RGB images with known poses and learning a 3D descriptor representation that models variations in appearance according to viewpoint. Our method is also suitable for integration with existing track-based localization systems for robotics.